HUBBLE SPACE TELESCOPE DAILY REPORT #827 PERIOD COVERED: 0800 EST Daily Status Report as of 050/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of PC 3603 (Snapshots of Protoplanetary Nebulae) The Planetary Camera was used to take a total of five images of the protoplanetary system PK249+01D1. The first image was reported to be blank, and the others were reported to be okay. 1.2 Completed GHRS/2 3862 (Excitation of Atmospheres of Planetary Satellites) Both the High Resolution Spectrograph and the Faint Object Spectrographs were used to make eight spectral observations of the moon Io while it enters and exits the shadow of Jupiter. All of the observations were received, and where possible displayed. The target acquisition were successful. The Rapid Mode spectra were normal. They were displayed for the principal investigator. 1.3 Completed PC 4140 (WF/PC Cal: Rapid Internal Monitor- Part 2) The Planetary Camera was used to obtain two internal flat observations which will be used to check for new forms of contamination, and the return of quantum efficiency hysteresis. Both images were received and displayed. Both internal flats were reported to look normal. 1.4 Completed GHRS/2 4065 (Dark Noise Monitoring HRS) The High Resolution Spectrograph was used to make one internal dark observations to monitor the detector for dark noise. The observation was received and displayed. The dark exposure was reported to look normal. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed GHRS/2 4792 (Hydrogen Lyman Alpha Emission Line Profiles) The High Resolution Spectrograph was used to make four observations of the line profiles of the Lyman Alpha dayglow on the central meridian of Jupiter at the equator and mid-latitudes. All four observations were received and displayed in real time. The first and thirds observations were wave length calibration images, and the others were spectra of the target. They all were reported to look okay. 1.7 Initiated GHRS/2 3372 (GHRS Point Spread Function Determination) The High Resolution Spectrograph was used to make point spread function observations of standard star HD141556. Four observations were received and displayed. The first and third observations were deflection calibrations and the other two were a spiral search and the peakup. The target was found in the ninth map. After the peakup the target flux was 45044. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 3372 (GHRS Point Spread Function Determination) 3.2 Complete FOS/RD/BL 4115 (Imaging and Spectroscopy of Nearby Quasars) 3.3 Complete WFC 1297 (GTO High Latitude Parallel Program) 3.4 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.5 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.6 Complete GHRS/2 3626 (Empirical Determination of Wind Velocity and Density Laws) 3.7 Complete GHRS/2 3706 (Nature of Gas Loops in Milky Way Halo) 3.8 Complete GHRS/2 4067 (GHRS 1st Order Grating Spectral Lamp Test) 3.9 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.10 Complete GHRS/2 4768 (Dynamics and Energy Balance of Interstellar Transition Regions 3.11 Complete PC 4347 (WF/PC calibration: Flood Retention Test) 3.12 Complete GHRS/2 4792 (Hydrogen Lyman Alpha Dayglow Emission Line Profiles) 3.13 Complete PC 4712 (WF/PC Calibration: Measles Monitor) 3.14 Complete GHRS/2 4766 (Boron in Population II Dwarfs, Part 2) 3.15 Complete GHRS/2 (Highly Ionized Nitrogen in Galactic Halo) 3.16 Complete FOC/96 (3591 (Mass Stars in Starburst Galaxies- FOC Images (Part 1)) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: The FGS #2 anomalous behavior is still under investigation. A series of tests to further ascertain the problem are in process. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-